At 10 am a gymnastics team began traveling to it's district finals meet 200 miles away. Team A traveled on the interstate at 65 mph and had a flat tire halfway to the meet. Team B traveled at 55 mph. Approximately how many minutes did Team A sit before Team B arrived to help ?
The book says the answer is 17 minutes but I keep getting 10 ?!?!?
: At 10 am a gymnastics team began traveling to it's district finals meet 200 miles away. Team A traveled on the interstate at 65 mph and had a flat tire halfway to the meet. Team B traveled at 55 mph. Approximately how many minutes did Team A sit before Team B arrived to help ?
The book says the answer is 17 minutes but I keep getting 10 ?!?!?

Answer by nerdybill(1129)

At 10 am a gymnastics team began traveling to it's district finals meet 200 miles away. Team A traveled on the interstate at 65 mph and had a flat tire halfway to the meet. Team B traveled at 55 mph. Approximately how many minutes did Team A sit before Team B arrived to help ?
.
Assuming both teams left at the same time, the "half-way" point is 100 miles.
.
Team A traveling at 65mph gets to the halfway point in:
1/65 * 100 = 100/65 = 1.583 hours
.
Team B traveling at 55mph gets to the halfway point in:
1/55 * 100 = 100/55 = 1.818 hours
.
The time Team A has to wait is:
1.818 - 1.583 = .279 hours
.
To convert to minutes:
.279 * 60 = 16.78
or approximately 17 minutes